20m : 70m = Rs 8 : Rs ______.
 20m : 70m = Rs 8 : Rs ______.

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

20m : 70m = Rs 8 : Rs 28.
Let us assume the missing number be P. Then, 20m : 70m =₹8 :₹P

20/70 = 8/P
P = (70 × 8)/20 P = 560/20
P = 56/2
P = 28

Therefore, 20m : 70m = Rs 8 : Rs 28.

Thanks